EXAMPLES OF HOPF ALGEBRAS OF DIMENSION 2^{m}

YEVGENIA KASHINA

Abstract. In this paper we construct two families of semisimple Hopf algebras of dimension 2^{n+1}, n ≥ 3. They are all constructed as Radford’s biproducts. For these examples and their duals we compute their grouplike elements, centers, character algebras and Grothendieck rings. Comparing these facts we are able to show that depending on the dimension, representatives of one of the families are selfdual.